OpenStudy (anonymous):

5^x^2=625^x. solve the equation

OpenStudy (anonymous):

I'm sorry but what does 5^x^2 mean?

OpenStudy (anonymous):

5 raise to the x raise to the second

OpenStudy (anonymous):

\[5^{x^2} = 625^x\] right?

OpenStudy (anonymous):

\[5^{x^2} = 625^x\]

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

take log of both sides and recall that:\[\log(a^b) = b*\log(a)\]

OpenStudy (anonymous):

i know the answer is 0.4, but i don't know why

OpenStudy (anonymous):

take the log of both sides like Euler said,

OpenStudy (anonymous):

\[\log(5^{x^2}) = \log(625^x)\]using rule:\[x^2 \log(5) = x \log(625)\] use calculator :)

OpenStudy (anonymous):

This isn't simplified x=2+3√,2−3√

OpenStudy (anonymous):

I'm building off of Euler comment.

OpenStudy (anonymous):

\[x[xlog(5) - \log(625)] = 0\]answer is x = 0 or \[xlog(5) - \log(625) = 0\] \[x = \frac{ \log625 }{ \log5 } = 4\]

OpenStudy (anonymous):

answer is 0, 4 [zero and four] and not 0.4 :)

OpenStudy (anonymous):

\[\text{Log}\left[5^{x^2}\right]\text{=}\text{Log}\left[625^x\right] \]\[x^2 \text{Log}[5]=4 x \text{Log}[5]\]\[x^2 -4 x =0\]\[(x-4) x=0 \]You can take it from here.